# On generating mapping class groups by pseudo-Anosov elements

### v2

@@ 1. Introduction
The problem of finding generating sets for groups is a classical one, and has been studied by many authors. The first finite generating set for $\mathcal{M}_{g}$ was given by Dehn [4], and the generating set consists of $2g(g-1)$ Dehn twists. After that, Lickorish [14] showed that $13g-$ 1 Dehn twists generate $\mathcal{M}_{g}$. Moreover, Humphries [7] proved that $12g+$ 1 Dehn twists suffice to generate $\mathcal{M}_{g}$ and that $2g$ (or less) Dehn twists cannot generate $\mathcal{M}_{g}$. It is well-known that $\mathcal{H}_{g}$ is generated by $12g+$ 1 Dehn twists (see [2]). If we do not require the generators of $\mathcal{M}_{g}$ to be Dehn twists then the number of generators can be reduced. In fact, Lu [15] showed that $\mathcal{M}_{g}$ is generated by three elements. Here, since $\mathcal{M}_{g}$ is not cyclic, at least two elements are needed to generate $\mathcal{M}_{g}$. For this, Wajnryb [27] gave a generating set for $\mathcal{M}_{g}$ consisting of two elements for $g\geq 1$, that is, the generating set is minimal. Since then, various minimal generating sets for $\mathcal{M}_{g}$ have been constructed (see, for example, [10, 29, 1]). The problem of finding torsion (or involution) generating sets has also been studied, for example, for finite simple groups. There have been many results on torsion (or involution) generating sets for $\mathcal{M}_{g}$ (see, for example, [17, 20, 16, 3, 9, 10, 19, 5, 12, 11, 28, 29]). In particular, Korkmaz gave a minimal torsion generating set of $\mathcal{M}_{g}$ consisting of two torsion elements of order $24g+$ 2 for $g\geq 3$ [10] (which was improved to order $g$ by Yildiz [29] for $g\geq 6$). A minimal generating set for $\mathcal{H}_{g}$ consisting of two torsion elements of order $24g+$ 2 and $22g+$ 2 was constructed by Korkmaz [10] for $g=1,2$ and Stukow [25] for $g\geq 3$. We note that $\mathcal{H}_{g}$ can not be generated by involutions for all $g\geq 1$ from homological reason by the work of Birman and Hilden [2].
@@
By Nielsen-Thurston classification, every element $f$ in $\mathcal{M}_{g}$ (resp. $\mathcal{H}_{g}$) has a representative homeomorphism $\phi$ which is of one of the following three types: (1) periodic i.e., $\phi^{m}=\mathrm{id}$ for some integer $m>0$, (2) reducible i.e., $\phi$ leaves invariant a finite collection of pairwise disjoint essential simple closed curves on $\Sigma_{g}$, or (3) pseudo-Anosov [6, 26]. An element $f$ in $\mathcal{M}_{g}$ is said to be periodic (resp. reducible or pseudo-Anosov) if a representative of $f$ is periodic (resp. reducible or pseudo-Anosov). Periodic elements in $\mathcal{M}_{g}$ are just torsion elements. Note that there is an element in $\mathcal{M}_{g}$ which is periodic and reducible as shown in Figure 2 in the next section. A simple example of a reducible but not periodic element is a Dehn twist, and a torsion element in $\mathcal{M}_{g}$ of order $24g+$ 2 is periodic but not reducible. On the other hand, if $f$ is a pseudo-Anosov element then it is neither periodic nor reducible.
@@
We next focus on minimal normal generating sets for $\mathcal{M}_{g}$. Since any Dehn twists about nonseparating simple closed curves are conjugate, the above-mentioned generating sets given by Lickorish [14] and Humphries [7] (resp. Birman-Hilden [2]) are normal generating sets of $\mathcal{M}_{g}$ (resp. $\mathcal{H}_{g}$). Moreover, the above-mentioned minimal torsion generating set in [10] (resp. [29]) consisting of two periodic but not reducible (resp. periodic and reducible) elements is also a normal one. With this background, Theorem 1.2 (resp. 1.3) below gives a minimal normal generating set for $\mathcal{M}_{g}$ consisting of two pseudo-Anosov (resp. reducible but not periodic) elements. In Theorem 1.4 blow, we give a minimal normal generating set for $\mathcal{H}_{g}$ consisting of two reducible but not periodic elements. However, we do not know whether $\mathcal{H}_{g}$ is generated by two conjugate periodic (resp. pseudo-Anosov) elements.
Another motivation for Theorem 1.2 comes from the work of Lanier and Margalit [13]. To explain this, we introduce some terminologies for pseudo-Anosov elements. Let $g\geq 2$. When $f$ is a pseudo-Anosov element in $\mathcal{M}_{g}$, $f$ has a representative $\phi$ such that there are transverse measured foliations $(\mathcal{F}^{s},\mu_{s})$ and $(\mathcal{F}^{u},\mu_{u})$ on $\Sigma_{g}$, and a real number $\lambda>1$ so that $\phi\cdot(\mathcal{F}^{s},\mu_{s})=(\mathcal{F}^{s},\lambda^{-1}\mu_{s})$ and $\phi\cdot(\mathcal{F}^{u},\mu_{u})=(\mathcal{F}^{u},\lambda\mu_{u})$. Then, $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$ are called the stable and unstable foliations, and $\lambda$ is called the dilatation or stretch factor of $\phi$. Note that two conjugate pseudo-Anosov elements have the same dilatation. With these terminologies, we can state Lanier and Margalit’s result. They showed that $\mathcal{M}_{g}$ is normally generated by pseudo-Anosov elements with arbitrarily large dilatations for each $g\geq 1$ (see, Theorem 1.3 and Proposition 4.2 in [13]). Theorem 1.2 is a refinement of this result in the sense that two such conjugate pseudo-Anosov elements suffice to generate $\mathcal{M}_{g}$ if $g\geq 9$.
@@
It is well-known that the mapping class group $\mathcal{M}_{g}$ of $\Sigma_{g}$ can be identified with the orbifold fundamental group of the moduli space $\mathbb{M}_{g}$ of Riemann surfaces of genus $g$. When we equiv $\mathbb{M}_{g}$ with the Teichmüller metric, pseudo-Anosov elements in $\mathcal{M}_{g}$ correspond to geodesic loops in $\mathbb{M}_{g}$. In particular, the logarithm of the dilatation of a pseudo-Anosov element in $\mathcal{M}_{g}$ is the length of a geodesic loop in $\mathbb{M}_{g}$. Then, we can rewrite Theorem 1.2 in terms of the geometry of normal covers of $\mathbb{M}_{g}$. Theorem 1.2 implies that we can take two (resp. three) geodesic loops in $\mathbb{M}_{g}$ with the same arbitrarily length that generate the orbifold fundamental group of $\mathbb{M}_{g}$ if $g\geq 9$ (resp. $g\geq 3$).
@@ 2. Preliminaries
  • If $a$ intersects $b$ transversely at exactly one point, then the braid relation $t_{a}t_{b}t_{a}=t_{b}t_{a}t_{b}$ holds. This gives $t_{b}t_{a}(b)=a$ (from $t_{t_{b}t_{a}(b)}=(t_{b}t_{a})t_{b}(t_{b}t_{a})^{-1}=t_{a}$), and hence $t_{a}(b)=t_{b}^{-1}(a)$.
  • Let $x,y,z$ be the interior curves on a subsurface of genus 00 with four boundary curves $a,b,c,d$ on $\Sigma_{g}$ as in Figure 1. Then, the lantern relation $t_{d}t_{c}t_{b}t_{a}=t_{z}t_{y}t_{x}$ holds. The lantern relation was discovered by Dehn [4] and rediscovered by Johnson [8].
We assume that the surface of this paper is the $yz$-plane and that $\Sigma_{g}$ is embedded in $\mathbb{R}^{3}$ as in Figure 2 (resp. Figure 3) such that it is invariant under the rotation $r$ by $-\frac{2\pi}{g}$ about the $x$-axis (resp. the rotation $\iota$ by $-\pi$ about the $y$-axis, i.e., the hyperelliptic involution, and the rotation $r_{\pi}$ by $-\pi$ about the $x$-axis). Note that $r_{\pi}$ is an element in $\mathcal{H}_{g}$ since $\iota r_{\pi}=r_{\pi}\iota$.
@@
The notations $a_{i},b_{i},c_{i}$ (resp. $\alpha_{j}$, $\beta$) will always denote the simple closed curves on $\Sigma_{g}$ shown in Figure 2 (resp. 3) for $i=1,2,\ldots,g$ (resp. $j=1,2,\ldots,2g+1$). Note that $\alpha_{1}=a_{1}$, $\alpha_{2i}=b_{i}$, $\alpha_{2i+1}=c_{i}$ for $i=1,2,\ldots,g-1$.
Figure 2. The rotation $r:\Sigma_{g}\to\Sigma_{g}$ by $-\frac{2\pi}{g}$ about the $x$-axis and the simple closed curves $a_{i},b_{i},c_{i}$ on $\Sigma_{g}$ for $i=1,2,\ldots,g$.
Figure 3. The hyperelliptic involution $\iota:\Sigma_{g}\to\Sigma_{g}$, the rotation $r_{\pi}:\Sigma_{g}\to\Sigma_{g}$ by $-\pi$ about the $x$-axis, and the simple closed curves $\alpha_{1},\alpha_{2},\ldots,\alpha_{2g+1},\beta$ on $\Sigma_{g}$.
@@
For $g\geq 1$, $\mathcal{M}_{g}$ is generated by $t_{a_{1}},\ldots,t_{a_{g}},t_{b_{1}},\ldots,t_{b_{g}}$ and $t_{c_{1}},\ldots,t_{c_{g-1}}$.
@@
For $g\geq 1$, $\mathcal{M}_{g}$ is generated by $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g}}$ and $t_{\beta}$.
Theorem 2.3 below is very useful to construct a pseudo-Anosov element. A multicurves in $\Sigma_{g}$ is the union of a finite collection of disjoint essential simple closed curves in $\Sigma_{g}$.
@@
$h_{2i-1}=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\cdot t_{\alpha_{2g+1}}\ \ \mathrm{and}$
$h_{2i}=t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdot t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\cdot t_{\alpha_{2g+1}}$
@@
$f_{2i-1}=t_{\beta}\cdot t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\ \ \mathrm{and}$
$f_{2i}=t_{\beta}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdot t_{\alpha_{2i+3}}t_{\alpha_{2i+4}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
Since $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are generators of $\mathcal{H}_{g}$ (see [2]), $h_{1}$ is also in $\mathcal{H}_{g}$. In the notation of Theorem 3.1 in [23], let $\mathscr{C}=\{\alpha_{1},\alpha_{3},\ldots,\alpha_{2g-1},\alpha_{2g+1}\}$ and $\mathscr{D}=\{\alpha_{2},\alpha_{4},\ldots,\alpha_{2g}\}$ be two multi-curves on $\Sigma_{g}$ (i.e., a set of pairwise disjoint simple closed curves). Then, these two multi-curves $\mathscr{C}$ and $\mathscr{D}$ satisfy the condition of Theorem 2.3, and therefore, we see that $h_{1}$ is a pseudo-Anosov element. By applying these arguments again, with $\mathscr{C}=\{\alpha_{1},\alpha_{3},\ldots,\alpha_{2g-1},\alpha_{2g+1}\}$ replaced by $\mathscr{C}^{\prime}=\{\alpha_{3},\alpha_{5},\ldots,\alpha_{2g+1}\}$, we see that $h_{2}$ is also a pseudo-Anosov in $\mathcal{H}_{g}$.
@@
If we take two multicurves $\mathscr{C}=\{a_{1},c_{1},a_{2},c_{2},\ldots,a_{g},c_{g}\}$ and $\mathscr{D}=\{b_{1},b_{2},\ldots,b_{g}\}$ then by Theorem 2.3, we see that $\rho_{n}^{g}$ is a pseudo-Anosov element. Therefore $\rho_{n}$ is a pseudo-Anosov.

### v3

@@ 1. Introduction
The problem of finding generating sets for groups is a classical one and has been studied by many authors. The first finite generating set for $\mathcal{M}_{g}$ was given by Dehn [4], and the generating set consists of 2g(g1)2g(g-1) Dehn twists. After that, Lickorish [14] showed that 3g13g-1 Dehn twists generate $\mathcal{M}_{g}$. Moreover, Humphries [7] proved that $2g+1$ Dehn twists suffice to generate $\mathcal{M}_{g}$ and that $2g$ (or less) Dehn twists cannot generate $\mathcal{M}_{g}$. It is well-known that $\mathcal{H}_{g}$ is generated by $2g+1$ Dehn twists (see [2]). If we do not require the generators of $\mathcal{M}_{g}$ to be Dehn twists, then the number of generators can be reduced. In fact, Lu [15] showed that $\mathcal{M}_{g}$ is generated by three elements. Here, since $\mathcal{M}_{g}$ is not cyclic, at least two elements are needed to generate $\mathcal{M}_{g}$. For this, Wajnryb [27] gave a generating set for $\mathcal{M}_{g}$ consisting of two elements for $g\geq 1$, that is, the generating set is minimal. Since then, various minimal generating sets for $\mathcal{M}_{g}$ have been constructed (see, for example, [10, 29, 1]). The problem of finding torsion (or involution) generating sets has also been studied, for example, for finite simple groups. There have been many results on torsion (or involution) generating sets for $\mathcal{M}_{g}$ (see, for example, [17, 20, 16, 3, 9, 10, 19, 5, 12, 11, 28, 29]). In particular, Korkmaz gave a minimal torsion generating set of $\mathcal{M}_{g}$ consisting of two torsion elements of order $4g+2$ for $g\geq 3$ [10] (which was improved to order $g$ by Yildiz [29] for $g\geq 6$). A minimal generating set for $\mathcal{H}_{g}$ consisting of two torsion elements of order $4g+2$ and $2g+2$ was constructed by Korkmaz [10] for $g=1,2$ and Stukow [25] for $g\geq 3$. We note that $\mathcal{H}_{g}$ can not be generated by involutions for all $g\geq 1$ from homological reasons by the work of Birman and Hilden [2].
@@
By Nielsen-Thurston classification, every element $f$ in $\mathcal{M}_{g}$ (resp. $\mathcal{H}_{g}$) has a representative homeomorphism $\phi$ which is of one of the following three types: (1) periodic i.e., $\phi^{m}=\mathrm{id}$ for some integer $m>0$, (2) reducible i.e., $\phi$ leaves invariant a finite collection of pairwise disjoint essential simple closed curves on $\Sigma_{g}$, or (3) pseudo-Anosov [6, 26]. An element $f$ in $\mathcal{M}_{g}$ is said to be periodic (resp. reducible or pseudo-Anosov) if a representative of $f$ is periodic (resp. reducible or pseudo-Anosov). Periodic elements in $\mathcal{M}_{g}$ are just torsion elements. Note that there is an element in $\mathcal{M}_{g}$ which is periodic and reducible as shown in Figure 2 in the next section. A simple example of a reducible but not periodic element is a Dehn twist, and a torsion element in $\mathcal{M}_{g}$ of order $4g+2$ is periodic but not reducible. On the other hand, if $f$ is a pseudo-Anosov element, then it is neither periodic nor reducible.
@@
We next focus on minimal normal generating sets for $\mathcal{M}_{g}$. Since any Dehn twists about nonseparating simple closed curves are conjugate, the above-mentioned generating sets given by Lickorish [14] and Humphries [7] (resp. Birman-Hilden [2]) are normal generating sets of $\mathcal{M}_{g}$ (resp. $\mathcal{H}_{g}$). Moreover, the above-mentioned minimal torsion generating set in [10] (resp. [29]) consisting of two periodic but not reducible (resp. periodic and reducible) elements is also a normal one. With this background, Theorem 1.2 (resp. 1.3) below gives a minimal normal generating set for $\mathcal{M}_{g}$ consisting of two pseudo-Anosov (resp. reducible but not periodic) elements. In Theorem 1.4 below, we give a minimal normal generating set for $\mathcal{H}_{g}$ consisting of two reducible but not periodic elements. However, we do not know whether $\mathcal{H}_{g}$ is generated by two conjugate periodic (resp. pseudo-Anosov) elements.
Another motivation for Theorem 1.2 comes from the work of Lanier and Margalit [13]. To explain this, we introduce some terminology for pseudo-Anosov elements. Let $g\geq 2$. When $f$ is a pseudo-Anosov element in $\mathcal{M}_{g}$, $f$ has a representative $\phi$ such that there are transverse measured foliations $(\mathcal{F}^{s},\mu_{s})$ and $(\mathcal{F}^{u},\mu_{u})$ on $\Sigma_{g}$, and a real number $\lambda>1$ so that $\phi\cdot(\mathcal{F}^{s},\mu_{s})=(\mathcal{F}^{s},\lambda^{-1}\mu_{s})$ and $\phi\cdot(\mathcal{F}^{u},\mu_{u})=(\mathcal{F}^{u},\lambda\mu_{u})$. Then, $\mathcal{F}^{s}$ and $\mathcal{F}^{u}$ are called the stable and unstable foliations, and $\lambda$ is called the dilatation or stretch factor of $\phi$. Note that two conjugate pseudo-Anosov elements have the same dilatation. With these terminologies, we can state Lanier and Margalit’s result. They showed that $\mathcal{M}_{g}$ is normally generated by pseudo-Anosov elements with arbitrarily large dilatations for each $g\geq 1$ (see Theorem 1.3 and Proposition 4.2 in [13]). Theorem 1.2 is a refinement of this result in the sense that two such conjugate pseudo-Anosov elements suffice to generate $\mathcal{M}_{g}$ if $g\geq 9$.
@@
It is well-known that the mapping class group $\mathcal{M}_{g}$ of $\Sigma_{g}$ can be identified with the orbifold fundamental group of the moduli space $\mathbb{M}_{g}$ of Riemann surfaces of genus $g$. When we equiv $\mathbb{M}_{g}$ with the Teichmüller metric, pseudo-Anosov elements in $\mathcal{M}_{g}$ correspond to geodesic loops in $\mathbb{M}_{g}$. In particular, the logarithm of the dilatation of a pseudo-Anosov element in $\mathcal{M}_{g}$ is the length of a geodesic loop in $\mathbb{M}_{g}$. Then, we can rewrite Theorem 1.2 in terms of the geometry of normal covers of $\mathbb{M}_{g}$. Theorem 1.2 implies that we can take two (resp. three) geodesic loops in $\mathbb{M}_{g}$ with the same arbitrary large length that generate the orbifold fundamental group of $\mathbb{M}_{g}$ if $g\geq 9$ (resp. $g\geq 3$).
@@ 2. Preliminaries
  • If $a$ intersects $b$ transversely at exactly one point, then $t_{b}t_{a}(b)=a$, and hence $t_{a}(b)=t_{b}^{-1}(a)$. This gives the braid relation $t_{a}t_{b}t_{a}=t_{b}t_{a}t_{b}$.
  • Let $x,y,z$ be the interior curves on a subsurface of genus 0 with four boundary curves $a,b,c,d$ on $\Sigma_{g}$ as in Figure 1. Then, the lantern relation $t_{d}t_{c}t_{b}t_{a}=t_{x}t_{y}t_{z}$ holds. The lantern relation was discovered by Dehn [4] and rediscovered by Johnson [8].
We assume that the surface of this paper is the $yz$-plane and that $\Sigma_{g}$ is embedded in $\mathbb{R}^{3}$ as in Figure 2 (resp. Figure 3) such that it is invariant under the rotation $r$ by $\frac{2\pi}{g}$ about the $x$-axis (resp. the rotation $\iota$ by $\pi$ about the $y$-axis, i.e., the hyperelliptic involution, and the rotation $r_{\pi}$ by $\pi$ about the $x$-axis). Note that $r_{\pi}$ is an element in $\mathcal{H}_{g}$ since $\iota r_{\pi}=r_{\pi}\iota$.
@@
The notations $a_{i},b_{i},c_{i}$ (resp. $\alpha_{j}$, $\beta$) will always denote the simple closed curves on $\Sigma_{g}$ shown in Figure 2 (resp. 3) for $i=1,2,\ldots,g$ (resp. $j=1,2,\ldots,2g+1$). We identify surfaces in Figure 2 and Figure 3 so that $\alpha_{1}=a_{1}$, $\beta=a_{2}$, $\alpha_{2i}=b_{i}$, $\alpha_{2i+1}=c_{i}$ for $i=1,2,\ldots,g-1$.
Figure 2. The rotation $r:\Sigma_{g}\to\Sigma_{g}$ by $\frac{2\pi}{g}$ about the $x$-axis and the simple closed curves $a_{i},b_{i},c_{i}$ on $\Sigma_{g}$ for $i=1,2,\ldots,g$.
Figure 3. The hyperelliptic involution $\iota:\Sigma_{g}\to\Sigma_{g}$, the rotation $r_{\pi}:\Sigma_{g}\to\Sigma_{g}$ by $\pi$ about the $x$-axis, and the simple closed curves $\alpha_{1},\alpha_{2},\ldots,\alpha_{2g+1},\beta$ on $\Sigma_{g}$.
@@
For $\geq 1, g\mathcal{M}_{g}$ is generated by $t_{a_{1}},\ldots,t_{a_{g}},t_{b_{1}},\ldots,t_{b_{g}}$ and $t_{c_{1}},\ldots,t_{c_{g-1}}$.
@@
For $\geq 1, g\mathcal{M}_{g}$ is generated by $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g}}$ and $t_{\beta}$.
Theorem 2.3 below is very useful to construct a pseudo-Anosov element. A multicurve in $\Sigma_{g}$ is the union of a finite collection of disjoint essential simple closed curves in $\Sigma_{g}$.
@@
$h_{2i-1}=t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\cdot t_{\alpha_{2g+1}},\ \ \mathrm{and}$
$h_{2i}=t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}\cdot t_{\alpha_{2g+1}}$
@@
$f_{2i-1}=t_{\beta}\cdot t_{\alpha_{2i-1}}t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1},\ \ \mathrm{and}$
$f_{2i}=t_{\beta}\cdot t_{\alpha_{2i}}^{-1}\cdot t_{\alpha_{2i+1}}t_{\alpha_{2i+2}}^{-1}\cdots t_{\alpha_{2g-1}}t_{\alpha_{2g}}^{-1}$
@@
Since $t_{\alpha_{1}},t_{\alpha_{2}},\ldots,t_{\alpha_{2g+1}}$ are generators of $\mathcal{H}_{g}$ (see [2]), $h_{1}$ is also in $\mathcal{H}_{g}$. In the notation of Theorem 3.1 in [23], let $\mathscr{C}=\{\alpha_{1},\alpha_{3},\ldots,\alpha_{2g-1},\alpha_{2g+1}\}$ and $\mathscr{D}=\{\alpha_{2},\alpha_{4},\ldots,\alpha_{2g}\}$ be two multicurves on $\Sigma_{g}$. Then, these two multicurves $\mathscr{C}$ and $\mathscr{D}$ satisfy the condition of Theorem 2.3, and therefore, we see that $h_{1}$ is a pseudo-Anosov element. By applying these arguments again, with $\mathscr{C}=\{\alpha_{1},\alpha_{3},\ldots,\alpha_{2g-1},$ $\alpha_{2g+1}\}$ replaced by $\mathscr{C}^{\prime}=\{\alpha_{3},\alpha_{5},\ldots,\alpha_{2g+1}\}$, we see that $h_{2}$ is also a pseudo-Anosov in $\mathcal{H}_{g}$.
@@
If we take two multicurves $\mathscr{C}=\{a_{1},c_{1},a_{2},c_{2},\ldots,a_{g},c_{g}\}$ and $\mathscr{D}=\{b_{1},b_{2},\ldots,b_{g}\}$ then by Theorem 2.3, we see that $\rho_{n}^{g}$ is a pseudo-Anosov element. Therefore, $\rho_{n}$ is a pseudo-Anosov.